Identify the element X in the following nuclear reactions.

(I) N^15 +H^1 ⟶ He^4 +X

(ii) H^1+Be^9 ⟶ X + n

(III) C^13 + H^1⟶ X


1
Expert's answer
2021-05-28T00:56:42-0400

Solution:

(I) The unknown element X must have a mass number of A = (15 + 1 - 4) = 12 and an atomic number of Z = (7 + 1 - 2) = 6. The element with Z = 6 is carbon (C), so the balanced nuclear equation is as follows:


Answer (I): C^12


(II) The unknown element X must have a mass number of A = (1 + 9 - 1) = 9 and an atomic number of Z = (1 + 4 - 0) = 5. The element with Z = 5 is boron (B), so the balanced nuclear equation is as follows:


Answer (II): B^9


(III) The unknown element X must have a mass number of A = (13 + 1) = 14 and an atomic number of Z = (6 + 1) = 7. The element with Z = 7 is nitrogen (N), so the balanced nuclear equation is as follows:


Answer (III): N^14